OpenStudy (alprincenofl):

Find the prime factorization of 111111

OpenStudy (anonymous):

evidently 3 divides it

OpenStudy (anonymous):

also 11

OpenStudy (anonymous):

you can go from there i am sure

OpenStudy (alprincenofl):

111111 = 3·7·11·13·37

OpenStudy (anonymous):

if you meant \[3×7×11×13×37\] then yes

OpenStudy (xxdragoonxx):

11111 / 41 = 271 11111 / 271 = 41

OpenStudy (anonymous):

wow all those question marks

OpenStudy (anonymous):

let me try again \[3\times 7\times 11\times 13\times 37\] then yes

OpenStudy (alprincenofl):

If you want to get the prime factorization of 111111, the first thing you should notice is that the number would be divisible by 11. 111111 / 11 = 10101 Then try smaller dividing by smaller numbers to break it down. 10101 / 3 = 3367 3367 / 7 = 481 481 / 13 = 37 So, when you get the last two factors to be both prime numbers, you do not have to go any further. Put all the numbers together and you can check your answer: 11 * 3 * 7 * 13 * 37 = 111111 The factors of 111111 are 11, 3, 7, 13, 37.